OpenStudy (anonymous):

What are the slope and y-intercept of the graph of the given equation? y = –4x + 2

OpenStudy (anonymous):

When it's in this form (slope intercept form) y = mx + b The slope is "m" and the y-intercept is "b"

OpenStudy (anonymous):

is this the answer ?? The slope is 4 and the y-intercept is –2.

OpenStudy (phi):

match y = –4x + 2 y = mx + b

OpenStudy (anonymous):

this one then ? The slope is –4 and the y-intercept is 2

OpenStudy (anonymous):

Yes, that's correct.

OpenStudy (anonymous):

thanks :)